OpenStudy (anonymous):

What is the surface area of a cone with a radius of 12 centimeters and a slant height of 8 centimeters? Use 3.14 for pi.

OpenStudy (anonymous):

formula is \[S=\pi rs + \pi r^2\] yes?

OpenStudy (anonymous):

502.4 cm^2 753.6 cm^2 803.84 cm^2 1,205.76cm^2

OpenStudy (anonymous):

in your case r = 8, s = 12

OpenStudy (anonymous):

Annnd which answer is that?

OpenStudy (anonymous):

dunno let me get a calculator and see what i get. do you have one?

OpenStudy (anonymous):

K

OpenStudy (anonymous):

oops i read it wrong r = 12 and s = 8 so you want \[\pi \times 8\times 12+\pi\times 12^2\]

OpenStudy (anonymous):

oh sorry...dunno where my mind is//very sleepy..sorry

OpenStudy (anonymous):

well do u have a ps3?

OpenStudy (anonymous):

what is with these numbers by the names?

OpenStudy (anonymous):

Idk, I copied and pasted

OpenStudy (anonymous):

ok i got 753.6 using 3.14 for pi